OpenStudy (anonymous):

okay i need help with a few geometry problems can someone help me through them, i dont just want answers i want to learnnn! <3

OpenStudy (anonymous):

Find the value of x. If your answer is not an integer, leave it in simplest radical form. The diagram is not drawn to scale

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (saifoo.khan):

Learn Pythagoras Theorm here: https://www.khanacademy.org/math/geometry/right_triangles_topic/pyth_theor/v/the-pythagorean-theorem

OpenStudy (anonymous):

using SOHCAHTOA, tan(theta)=opp/adj tan(60)=(x/5),x=5tan(60), x=8.66 or Pythagoras's theorem x^2+y^2=z^2, 5^2+x^2=10^2, 25+x^2=100, x^2=100-25, x=sqrt(75)= 8.66

OpenStudy (anonymous):

holy crap lol, okay lets do this using pythag theorem.. step by step now thats too much for me lol. a2+b2=c2 so 5^2+ x^2 = 10^2 right?

OpenStudy (saifoo.khan):

right.

OpenStudy (saifoo.khan):

Solve for x.

OpenStudy (anonymous):

.. idk how to do that part.. 10^2-5^2?

OpenStudy (saifoo.khan):